Grade 12th passMost Scoring Topics in IIT JEE

Two stars of masses m and 2m are co-rotating about there centre of mass. Their centres are at a distance R apart. If R is much larger than the sizes of the stars , their common period of revolution is proportional to

RITESH
8 Years agoGrade 12th pass

2 Answers

TECH HACK
8 Years ago
Is it directly proportional to r^3/2If yes then Solution:F=GM 2m/r^2And F=mV^2/r[V is velocity]So V = root(2Gm/r^3)And V = r omegaAnd T=2π/omega Therefore T = 2πr√(r^3)/√()
TECH HACK
8 Years ago
Sorry the answer It was left uncompleted.Continue the solution from hereT = 2π√(r^3)/√(GM)Therefore T is directly proportional to √(r^3)